When Milo got promoted at work, he received a 25% pay raise. He now earns $60,000 per year.
What was his annual salary before his raise?

Answer by ikleyn(52776)

.

    x + 0.25x = 60000  dollars

      1.25x   = 60000

          x   = 60000/1.25 = 48000.    ANSWER

Solved.

The equations are self-explanatory.